OpenStudy (anonymous):

Easy stuff: Sum of 3 consecutive integers is 53 more than the smallest integer. Find em.

OpenStudy (anonymous):

x+x+1+x+2=53+x 3x + 3 = 53 +x 2x = 50 x =25 25,26,27 check - 25 + 26 + 27 = 78 53+25 = 78 yeah very easy stuff.

OpenStudy (anonymous):

Juanita I already got it but I can still give u a medal :p

OpenStudy (anonymous):

HAHA...well the question was still open :)